The starter motor of a car engine draws a current of 110 A from the battery. The copper wire to the motor is 4.00 mm in diameter and 1.2 m long. The starter motor runs for 0.920 s until the car engine starts.

How far does an electron travel along the wire while the starter motor is on? in mm.
Answer #1


I =110 A, L = 1.2 m , d = 4 mm ,r = 2 mm, t = 0.92 s

current denisty J = I/A

J = 110/(3.14*0.002^2) = 8.76*10^6 A/m

drift velocity v = J/ne

copper have n = 8.46*10^28 electrons per cubic meter

v = 8.76*10^6/(8.46*10^28*1.6*10^-19)

v = 647*10^-6 m/s

x = vt = 647*10^-6*0.92

x = 0.595 mm